Grade 11Physical Chemistry

two oxides of metal M have27.6% and 30% oxygen by weight . if the formula of the first oxide is M3O4.What is the formula of second oxide

Bhavya G
8 Years agoGrade 11

1 Answer

Jobin C J
8 Years ago
In the First Oxide Oxygen= 27.6% ; ∴Metal=100-27.6=72.4 Parts ;In the Second Oxide Oxygen= 30% ; ∴Metal=100-30=70 Parts ;As First Oxide Is M₃O₄ ;72.4 Part of M=3 Atoms of M ;27.6 Parts of O=4 Atoms of Oxygen;∴70 Parts of M=(3/72.4)×70 atoms of M = 2.9 atoms of M ;30 Parts of O=(4/27.6)×30 atoms of O = 4.35 atoms of O ;∴Ratio of M:O in the 2nd Oxide = 2.9 : 4.35 = 1 : 1.5 = 2 : 3 ;Hence Formula = M₂O₃